Shift Summary, Aug. 7, 1999 Day Shift

Richard Ford (rickford@fnal.gov)
Sat, 7 Aug 1999 15:45:44 -0500 (CDT)

Shift Summary for DAY shift, 07/AUG/99 (Director's cut)

On Shift: RF, VO

Hours of beam available at all intensities: 4.5/8.5
Hours of beam available at nominal intensity: 4.5/8.5
Hours of E832 data (special runs not included): 3.5/8.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: plane 4 crashes

========================================================================
Runs ended this shift: None

Tapes brought to FCC this shift: 228-230,080
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:

E832 datataking, Beam is very intermittent with lots of aborts.
Tev injection problems

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14134 07:30 Stream 6 DC2 crashed, reboot failed, power cycled
DRT_FRESH_START. Start run 14135

14135 45 8:55 Stream 6 plane 4 DC2 crashed again, reseated the cards,
changed tapes,DRT_FRESH_START

14137 9:28 Start Run 14137

21 9:56 KFREND

10:05 Shift crew goes to make coffee. No coffee. Crew gets
edgy.

60 10:45 Beam dies.

11:15 One of the shift crew members makes continuous
references to some guy named Barton from Atlanta.
Other member remains edgy.

98 11:40 Beam back.

11:50 Run manager paged to discuss coffee situation.
Run manager claims spokesperson responsible for
coffee delivery.

124 12:15 Beam goes away.

132 12:25 MCR calls to tell us that they are having injection
problems. A TeV expert is in working on it. No est.

12:30 Jason Hamm wanders in and points out that
we have 2 onsplit tape mount requests. We
immediately jump to it and mount the tapes. It's
amazing what you miss when you have no coffee.

145 12:40 Beam back. SEM very high.

12:45 SEM way too high (7.8E12) for one spill.
Call MCR and have them tone it down.

157 12:55 Beam now with reasonable SEM and centered.
Spills 145-156 have crappy beam.

158 12:58 Beam going away again...

13:10 No coffee has appeared. More references to Barton.
Other shift crew member offers doughnuts in a
conciliatory manner. Offer is greeted with
derision.

13:15 Shift crew member 1, recognizing survival depends
crucially on rapid coffee replenishment, offers
to go get coffee. Shift crew member 2 growls and
locks door. Both members remain edgy.

13:20 Shift crew member 2 begins referring to the new
California assault rifle ban as stupid and unnecessary.

183 13:31 What little beam we were getting has disappeared.

13:34 Spokesperson arrives with coffee. Just in time.

188 13:37 Beam back...

211 14:10 Etotal was inefficient. (Stripe in the HCC illumination
plot). Jason reloads Et. What a guy!

14:30 Notice the JFK face on one of the KQ trigger readout
plots. Page Jim, who is coming in.

15:00 Jim comes in and makes a frowny face. Calls Jason
and Peter at a scientific conference and leaves
message. The problem is only in the readout so
everything is probably ok until Tuesday when a rogue
blossom board will be replaced.

15:00 Beam gone...

========================================================================
Follow up Notes on Previous Shift

Although we started the new L3 processes, HCC still occasionally
gives false alarms.

========================================================================
Items to Follow up on Next Shift

If beam disappears for more than an hour, please run a laser scan.

Rid hall of extraneous assault rifles.